# Coding for Gaussian Two-Way Channels: Linear and Learning-Based Approaches

### v1

@@ III-C Conversion of Optimization Problem
The optimal value of ${\max}\big{\{}\mathbb{E}\big{[}\|{\bf x}_{1}\|^{2}\big{]},\mathbb{E}\big{[}\|{\bf x}_{2}\|^{2}\big{]}\big{\}}$ in (28) is an increasing function of $\eta_{1}$ when $\eta_{2}$ is fixed.
The proof of Lemma 1 is provided in Appendix C. The above lemma allows us to develop an efficient search method along $\eta_{1}$, which will be discussed in detail in Sec. III-F. Nevertheless, solving (28) is still intractable. As an approximation technique, we use the weighted sum method, which is a common choice in multi-objective optimization [21]. This method involves solving a weighted sum of the objective functions, and then varying the weights to generate a set of solutions that lie on the Pareto front. The optimal solution is then selected from this set based on the original objective function. In our case, we select the solution that minimizes ${\max}\big{\{}\mathbb{E}\big{[}\|{\bf x}_{1}\|^{2}\big{]},\mathbb{E}\big{[}\|{\bf x}_{2}\|^{2}\big{]}\big{\}}$.
Based on the weighted sum method, we introduce a weighting coefficient $\alpha\in(0,1)$ and aim to minimize the weighted sum of the users’ transmit powers under their SNR constraints. The corresponding optimization problem is given by
$\underset{{\bf g}_{1},{\bf F}_{1},{\bf g}_{2},{\bf F}_{2}}{\text{minimize}}$ $\alpha\mathbb{E}\big{[}\|{\bf x}_{1}\|^{2}\big{]}+(1-\alpha)\mathbb{E}\big{[}\|{\bf x}_{2}\|^{2}\big{]}$
@@
In summary, to achieve equivalence between the original sum-error minimization problem in (25) and the weighted sum-power minimization problem in (29) in terms of obtaining optimal solutions [21, 22], we solve (29) for every feasible solution sets of $(\eta_{1},\eta_{2},\alpha)$ and find the best solutions of ${\bf g}_{1}$, ${\bf F}_{1}$, ${\bf g}_{2}$, and ${\bf F}_{2}$, that minimize $\mathcal{E}_{1}(\text{SNR}_{1})+\mathcal{E}_{2}(\text{SNR}_{2})$ while satisfying $\mathbb{E}\big{[}\|{\bf x}_{1}\|^{2}\big{]}\leq NP$ and $\mathbb{E}\big{[}\|{\bf x}_{2}\|^{2}\big{]}\leq NP$. Since the problem in (29) is more tractable, we first focus on solving it in Sec. III-D and III-E. Then, in Sec. III-F, we provide an efficient algorithm to find the solutions for our original problem in (25).
@@ III-D Optimization Problem for Weighted Sum-Power Minimization
For convenience, we define a set $\mathcal{F}_{2}$ as a solution space for ${\bf F}_{2}$, i.e., ${\bf F}_{2}\in\mathcal{F}_{2}$.
First, we will investigate the solution behavior for feedback of User 2. Specifically, we reveal in the following proposition that it is optimal for User 2 not to utilize the last channel use for feedback to User 1, i.e., $f_{2,N}=0$, regardless of the encoding schemes.
@@
The proof of Proposition 1 is provided in Appendix D. We next look into the solution behavior of the message encoding vector for User 2, ${\bf g}_{2}$. To this end, we first formulate the optimization problem (29) only with respect to ${\bf g}_{2}$ by using the expression in (16)-(17) as follows:
@@
We first conduct a matrix decomposition on ${\bf Q}_{2}$ and obtain ${\bf Q}_{2}=({\bf Q}_{2}^{1/2})^{2}$ where ${\bf Q}_{2}^{1/2}=({\bf Q}_{2}^{1/2})^{\top}$. [^2] If we conduct the singular value decomposition on ${\bf Q}_{2}$, we have ${\bf Q}_{2}={\bf U}{\Sigma}{\bf U}^{\top}$ and obtain ${\bf Q}_{2}^{1/2}={\bf U}{\Sigma}^{1/2}{\bf U}^{\top}$. Defining ${\bf q}_{2}={\bf Q}_{2}^{-1/2}{\bf g}_{2}$, we can obtain an equivalent optimization problem as
@@
${\bf B}=\alpha{\bf Q}_{2}^{1/2}{\bf F}_{1}^{\top}{\bf F}_{1}{\bf Q}_{2}^{1/2}+(1-\alpha){\bf Q}_{2}^{1/2}({\bf I}+{\bf F}_{2}{\bf F}_{1})^{\top}({\bf I}+{\bf F}_{2}{\bf F}_{1}){\bf Q}_{2}^{1/2}.$ (33)
We then introduce our conjecture on the objective function value of (32), based on which we find the optimal solution for ${\bf g}_{2}$ in (31).
@@
For any ${\bf F}_{1}$ and ${\bf F}_{2}\in\mathcal{F}_{2}$,
$\min\{\alpha\sigma_{1}^{2},(1-\alpha)\sigma_{2}^{2}\}\leq\nu_{\min}[{\bf B}]\leq(1-\alpha)\sigma_{2}^{2},$ (34)
where $\nu_{\min}[{\bf B}]$ denotes the smallest eigenvalue of ${\bf B}$ in (32).
We have proved a special case of this conjecture when $N=3$ in Appendix E. We note that, for any $N$, we have not found any example that violates the conjecture in our extensive numerical simulations, where ${\bf F}_{1}$ and ${\bf F}_{2}\in\mathcal{F}_{2}$ are randomly generated.
@@
If Conjecture 1 is true, ${\bf g}_{2}=[0,...,0,\sqrt{\eta_{2}}\sigma_{2}]^{\top}$ is optimal in (29) when $\alpha\geq\frac{\sigma_{2}^{2}}{\sigma_{1}^{2}+\sigma_{2}^{2}}$.
The proof of Proposition 2 is provided in Appendix F. The result of Proposition 2 shows that it is optimal for User 2 to transmit the message only over the last channel use when the weight coefficient in (29) satisfies $\alpha\geq\frac{\sigma_{2}^{2}}{\sigma_{1}^{2}+\sigma_{2}^{2}}$.
We can extract further implications from Proposition 2. In the context of feedback utilization [8, 10, 9, 6], one user transmits its message initially and exchanges symbols containing the message and feedback information with the other user. However, given that User 2 only transmits its message in the last channel use from Proposition 2, User 2 does not utilize feedback. Consequently, User 1 does not function as a helper in providing feedback to User 2. Therefore, we can state the following remark.
@@
Using Propositions 1 and 2, we next aim to simplify our optimization problem in (29). In our optimization, we consider $\alpha\geq\frac{\sigma_{2}^{2}}{\sigma_{1}^{2}+\sigma_{2}^{2}}$. From Proposition 2, we have ${\bf g}_{2}=[0,...,0,\sqrt{\eta_{2}}\sigma_{2}]^{\top}$ as an optimal solution, which always satisfies $\text{SNR}_{2}=\eta_{2}$ regardless of other variables. Thus, we can remove the dependency of the constraint for $\text{SNR}_{2}$ in (29). Further, to make (29) more tractable, we first perform a matrix decomposition on ${\bf Q}_{1}$ and obtain ${\bf Q}_{1}=({\bf Q}_{1}^{1/2})^{2}$ where ${\bf Q}_{1}^{1/2}=({\bf Q}_{1}^{1/2})^{\top}$ in the same way for the decomposition on ${\bf Q}_{2}$ as described below (31). Then, we define ${\bf q}_{1}={\bf Q}_{1}^{-1/2}{\bf g}_{1}$ where ${\bf Q}_{1}=({\bf Q}_{1}^{1/2})^{2}$, which implies $\text{SNR}_{1}=\|{\bf q}_{1}\|^{2}$ and $\|{\bf g}_{1}\|^{2}={\bf q}^{\top}_{1}{\bf Q}_{1}{\bf q}_{1}$. Consequently, we rewrite the transmit powers in (16) and (17) as
$\mathbb{E}[\|{\bf x}_{1}\|^{2}]$ $={\bf q}^{\top}_{1}{\bf Q}_{1}{\bf q}_{1}+\|{\bf F}_{1}{\bf F}_{2}\|_{F}^{2}\sigma_{1}^{2}+\|{\bf F}_{1}\|_{F}^{2}\sigma_{2}^{2}$
@@
$\mathbb{E}[\|{\bf x}_{2}\|^{2}]$ $=\|{\bf g}_{2}\|^{2}+\|{\bf F}_{2}{\bf Q}_{1}^{1/2}{\bf q}_{1}\|^{2}+\|{\bf F}_{2}({\bf I}+{\bf F}_{1}{\bf F}_{2})\|_{F}^{2}\sigma_{1}^{2}+\|{\bf F}_{2}{\bf F}_{1}\|_{F}^{2}\sigma_{2}^{2},$ (36)
@@
$\underset{{\bf q}_{1},{\bf F}_{1},{\bf F}_{2}\in\mathcal{F}_{2}}{\text{minimize}}$ $\alpha\mathbb{E}\big{[}\|{\bf x}_{1}\|^{2}\big{]}+(1-\alpha)\mathbb{E}\big{[}\|{\bf x}_{2}\|^{2}\big{]}$

### v2

@@ III-C Conversion of Optimization Problem
The optimal solutions of (41) can be obtained by solving the following problem, defined as
subject to $({\bf g}_{1},{\bf F}_{1},{\bf g}_{2},{\bf F}_{2})\in\mathcal{S},$ (45)
where $\mathcal{S}=\bigcup_{\eta_{1},\eta_{2}\in\mathbb{R}^{+}}\mathcal{S}_{(\eta_{1},\eta_{2})}$. Here, $\mathcal{S}_{(\eta_{1},\eta_{2})}$ is the set of solutions $({\bf g}_{1},{\bf F}_{1},{\bf g}_{2},{\bf F}_{2})$ of (44) given $\eta_{1}$ and $\eta_{2}$, while satisfying ${\max}\big{\{}\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]}$, $\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{2}\|^{2}\big{]}\big{\}}$ $\leq NP$.
The proof of Lemma 1 is provided in Appendix E. Using Lemma 1, instead of directly tackling (41), we solve (45). Specifically, we first determine the feasible solution set $\mathcal{S}$ by solving (44) for all possible values of $\eta_{1}$ and $\eta_{2}$. After identifying $\mathcal{S}$, we find the solution that minimizes the sum-error. It is important to note that identifying $\mathcal{S}$ is the key to solve (45). Next, we derive the following lemma, which is useful to find $\mathcal{S}$.
Lemma 2.
The optimal value of ${\max}\big{\{}\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]},\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{2}\|^{2}\big{]}\big{\}}$ in (44) is an increasing function of $\eta_{1}$ when $\eta_{2}$ is fixed.
The proof of Lemma 2 is provided in Appendix F. Given that the search space for $\eta_{1}$ and $\eta_{2}$ is large in finding $\mathcal{S}$, we can use Lemma 2 to reduce the search space of $\eta_{1}$ once $\eta_{2}$ is fixed. That is, Lemma 2 allows us to develop an efficient search method along $\eta_{1}$, which will be discussed in detail in Sec. III-F.
To provide further tractability in solving (44), which is a step for finding $\mathcal{S}$ in (45), we treat the min-max problem in (44) as a multi-objective problem, where we aim to minimize both $\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]}$ and $\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{2}\|^{2}\big{]}$ and find the solution that minimizes ${\max}\big{\{}\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]},\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{2}\|^{2}\big{]}\big{\}}$. To solve the multi-objective problem, we use the weighted sum method as an approximated solution technique, which is a common choice in solving multi-objective problem [21]. We note that, using the weighted sum method provides benefits in terms of tractability, in our case for developing linear coding methods. We formulate the weighted sum-power minimization problem with a weighting coefficient $\alpha\in(0,1)$ as
$\underset{{\bf g}_{1},{\bf F}_{1},{\bf g}_{2},{\bf F}_{2}}{\text{minimize}}$ $\alpha\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]}+(1-\alpha)\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{2}\|^{2}\big{]}$
@@
The solution obtained from solving (46) for a specific weight $\alpha$ is feasible for (44), but it may not necessarily be optimal for (44). In general, solving weighted sum minimization problems does not guarantee to yield optimal solutions to non-convex multi-objective problems [22]. We note that choosing a proper weight in (46) is crucial for finding the optimal or sub-optimal solution to (44). However, it is challenging to determine the appropriate weight to achieve the optimal or sub-optimal solution of the min-max problem [22]. A straightforward approach is to (i) solve (46) given various values of weight $\alpha\in(0,1)$ to obtain feasible solutions for (44) and (ii) determine the solution that minimizes the max-power among the feasible solutions. We develop an efficient search method along $\alpha$, which will be discussed in detail in Sec. III-F.
In summary, we transform the optimization problem from solving the sum-error minimization problem in (41) to (45). The equivalence in optimization conversion is demonstrated through Lemma 1. Our focus then shifts to identifying the feasible solution set $\mathcal{S}$. Finding $\mathcal{S}$ involves solving the max-power minimization problem in (44) by varying the variables $\eta_{1}$ and $\eta_{2}$. To make it more tractable to find $\mathcal{S}$, we approximate solving (44) to solving the weighted-sum power minimization problem in (46) by varying the additional variable $\alpha$. While the optimization transform from (41) to (45) preserves optimality, we may lose some optimality when finding $\mathcal{S}$ by solving (46) instead of (44). The optimality is sacrificed for tractability during this step.
Solving (45) involves the following steps. First, we identify the feasible solution set $\mathcal{S}$ by solving (46) for various combinations of $(\eta_{1},\eta_{2},\alpha)$. Then, we determine the solution that minimizes the sum-error within $\mathcal{S}$. As the initial step in developing solution methods, we focus on solving (46) for a given $(\eta_{1},\eta_{2},\alpha)$ in Sec. III-D and III-E. Subsequently, in Sec. III-F, we provide an efficient algorithm to identify the feasible solution set $\mathcal{S}$ by varying $(\eta_{1},\eta_{2},\alpha)$, and ultimately, to find the solution for (45).
@@ III-D Optimization Problem for Weighted Sum-Power Minimization
For convenience, we define a set $\mathcal{F}_{2}$ as a solution space for ${\bf F}_{2}$, i.e., ${\bf F}_{2}\in\mathcal{F}_{2}$. First, we will investigate the solution behavior for feedback of User 2. Specifically, we reveal in the following proposition that it is optimal for User 2 not to utilize the last channel use for feedback to User 1, i.e., $f_{2,N}=0$, regardless of the encoding schemes.
@@
The proof of Proposition 1 is provided in Appendix G. We next look into the solution behavior of the message encoding vector for User 2, ${\bf g}_{2}$. To this end, we first formulate the optimization problem (46) only with respect to ${\bf g}_{2}$ by using the expression in (16)-(17) as follows:
@@
We note that ${\bf Q}_{2}$ is symmetric, so we can express it as ${\bf Q}_{2}={\bf U}{\Sigma}{\bf U}^{\top}$ by using the singular value decomposition. By defining ${\bf Q}_{2,\text{sqrt}}={\bf U}{\Sigma}_{\text{sqrt}}{\bf U}^{\top}$, where ${\Sigma}_{\text{sqrt}}$ is the diagonal matrix with entries that are the square roots of the corresponding diagonal entries in ${\Sigma}$, we rewrite ${\bf Q}_{2}$ as ${\bf Q}_{2}={\bf Q}_{2,\text{sqrt}}^{\top}{\bf Q}_{2,\text{sqrt}}$. Note that ${\bf Q}_{2,\text{sqrt}}={\bf Q}_{2,\text{sqrt}}^{\top}$. Since ${\bf Q}_{2}$ is invertible, ${\bf Q}_{2,\text{sqrt}}$ is also invertible. Defining ${\bf q}_{2}={\bf Q}_{2,\text{sqrt}}^{-1}{\bf g}_{2}$, we can obtain an equivalent optimization problem as
@@
${\bf B}=\alpha{\bf Q}_{2,\text{sqrt}}{\bf F}_{1}^{\top}{\bf F}_{1}{\bf Q}_{2,\text{sqrt}}+(1-\alpha){\bf Q}_{2,\text{sqrt}}({\bf I}+{\bf F}_{2}{\bf F}_{1})^{\top}({\bf I}+{\bf F}_{2}{\bf F}_{1}){\bf Q}_{2,\text{sqrt}}.$ (50)
We provide the following proposition to understand the solution behavior of ${\bf g}_{2}$ for any strictly lower triangular matrix ${\bf F}_{1}$ and any ${\bf F}_{2}\in\mathcal{F}_{2}$.
Proposition 2.
For any strictly lower triangular matrix ${\bf F}_{1}$ and any ${\bf F}_{2}\in\mathcal{F}_{2}$,
$0\leq\nu_{\min}[{\bf B}]\leq(1-\alpha)\sigma_{2}^{2},$ (51)
where $\nu_{\min}[{\bf B}]$ denotes the smallest eigenvalue of ${\bf B}$ in (50).
The proof of Proposition 2 is provided in Appendix H. We note that the value of $\nu_{\min}[{\bf B}]$ depends on the specific configurations of ${\bf F}_{1}$ and ${\bf F}_{2}$. We find that the lower bound of 0 is loose in Proposition 2. Intuitively, it is unlikely that $\nu_{\min}[{\bf B}]=0$, because this would imply that a certain choice of ${\bf q}_{2}$ (equivalently, a message encoding vector ${\bf g}_{2}$) could achieve zero transmit power (i.e., an objective value of 0) while still satisfying its message SNR constraint in (49). We recognize that there may be a tighter lower bound. Therefore, we introduce the following conjecture on the tighter lower bound of $\nu_{\min}[{\bf B}]$.
@@
For any strictly lower triangular matrix ${\bf F}_{1}$ and any ${\bf F}_{2}\in\mathcal{F}_{2}$,
$\min\{\alpha\sigma_{1}^{2},(1-\alpha)\sigma_{2}^{2}\}\leq\nu_{\min}[{\bf B}].$ (52)
We have proved a special case of this conjecture when $N=3$ in Appendix I. We note that, for any $N$, we have not found any example that violates the conjecture in our extensive numerical simulations, where ${\bf F}_{1}$ and ${\bf F}_{2}\in\mathcal{F}_{2}$ are randomly generated.
@@
If Conjecture 1 is true, ${\bf g}_{2}=[0,...,0,\sqrt{\eta_{2}}\sigma_{2}]^{\top}$ is optimal in (46) for any strictly lower triangular ${\bf F}_{1}$ and any ${\bf F}_{2}\in\mathcal{F}_{2}$ when $\alpha\geq\frac{\sigma_{2}^{2}}{\sigma_{1}^{2}+\sigma_{2}^{2}}$. This choice of ${\bf g}_{2}$ is optimal for any $\alpha\in[0,1]$ when $N=3$.
The proof of Proposition 3 is provided in Appendix J. The result of Proposition 3 shows that it is optimal for User 2 to transmit the message only over the last channel use when the weight coefficient in (46) satisfies $\alpha\geq\frac{\sigma_{2}^{2}}{\sigma_{1}^{2}+\sigma_{2}^{2}}$.
Corollary 1.
If Conjecture 1 is true, when $\alpha<\frac{\sigma_{2}^{2}}{\sigma_{1}^{2}+\sigma_{2}^{2}}$, the optimality loss of the objective value in (46) due to the choice of ${\bf g}_{2}=[0,...,0,\sqrt{\eta_{2}}\sigma_{2}]^{\top}$ in the worst case is $\eta_{2}(\sigma_{1}^{2}+\sigma_{2}^{2})(\frac{\sigma_{2}^{2}}{\sigma_{1}^{2}+\sigma_{2}^{2}}-\alpha)$.
The proof of Corollary 1 is provided in Appendix K. [^2] If Conjecture 1 is not true, the optimality loss in (46) due to the choice of ${\bf g}_{2}=[0,...,0,\sqrt{\eta_{2}}\sigma_{2}]^{\top}$ in the worst case is $(1-\alpha)\sigma_{2}^{2}\eta_{2}$. The proof follows almost identically to that in Appendix K, with the key difference being that the lower bound of $\nu_{\min}[{\bf B}]$ is zero from Proposition 2. We note from Proposition 3 that there is no optimality loss when $\alpha\geq\frac{\sigma_{2}^{2}}{\sigma_{1}^{2}+\sigma_{2}^{2}}$. We can extract further implications from Proposition 3. In the context of feedback utilization [8, 10, 9, 6], one user transmits its message initially and exchanges symbols containing the message and feedback information with the other user. However, given that User 2 only transmits its message in the last channel use from Proposition 3, User 2 does not utilize feedback. Consequently, User 1 does not function as a helper in providing feedback to User 2. Therefore, we can state the following remark.
@@
Using Propositions 1 and 3, we next aim to simplify our optimization problem in (46). In our optimization, we consider $\alpha\geq\frac{\sigma_{2}^{2}}{\sigma_{1}^{2}+\sigma_{2}^{2}}$. From Proposition 3, we have ${\bf g}_{2}=[0,...,0,\sqrt{\eta_{2}}\sigma_{2}]^{\top}$ as an optimal solution, which always satisfies $\text{SNR}_{2}=\eta_{2}$ regardless of other variables. Thus, we can remove the dependency of the constraint for $\text{SNR}_{2}$ in (46). Further, to make (46) more tractable, we first perform a matrix decomposition on ${\bf Q}_{1}$ and obtain ${\bf Q}_{1}={\bf Q}_{1,\text{sqrt}}^{\top}{\bf Q}_{1,\text{sqrt}}$ in the same way for the decomposition on ${\bf Q}_{2}$ as described below (48). Note that ${\bf Q}_{1,\text{sqrt}}={\bf Q}_{1,\text{sqrt}}^{\top}$. Since ${\bf Q}_{1}$ is invertible, ${\bf Q}_{1,\text{sqrt}}$ is also invertible. Then, we define ${\bf q}_{1}={\bf Q}_{1,\text{sqrt}}^{-1}{\bf g}_{1}$, which implies $\text{SNR}_{1}=\|{\bf q}_{1}\|^{2}$ and $\|{\bf g}_{1}\|^{2}={\bf q}^{\top}_{1}{\bf Q}_{1}{\bf q}_{1}$. Consequently, we rewrite the transmit powers in (16) and (17) as
$\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]}$ $={\bf q}^{\top}_{1}{\bf Q}_{1}{\bf q}_{1}+\|{\bf F}_{1}{\bf F}_{2}\|_{F}^{2}\sigma_{1}^{2}+\|{\bf F}_{1}\|_{F}^{2}\sigma_{2}^{2}$
@@
$\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{2}\|^{2}\big{]}$ $=\|{\bf g}_{2}\|^{2}+\|{\bf F}_{2}{\bf Q}_{1,\text{sqrt}}{\bf q}_{1}\|^{2}+\|{\bf F}_{2}({\bf I}+{\bf F}_{1}{\bf F}_{2})\|_{F}^{2}\sigma_{1}^{2}+\|{\bf F}_{2}{\bf F}_{1}\|_{F}^{2}\sigma_{2}^{2},$ (54)
@@
$\underset{{\bf q}_{1},{\bf F}_{1},{\bf F}_{2}\in\mathcal{F}_{2}}{\text{minimize}}$ $\alpha\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{1}\|^{2}\big{]}+(1-\alpha)\mathbb{E}_{{{{\color[rgb]{0,0,0}m_{1},m_{2},{\bf n}_{1},{\bf n}_{2}}}}}\big{[}\|{\bf x}_{2}\|^{2}\big{]}$